# FUSION: Forecast-Embedded Agent Scheduling with Service Incentive Optimization over Distributed Air-Ground Edge Networks

### v1

@@ Appendix C Proof of Property of Off-AIC2
$\bullet$ Critical reward (Phase 3). In Phase 3, for each winning AgP $v_{k}$ (with $\sum_{i}x^{\mathrm{Off}}_{i,k}\geq 1$), Off-AIC2 treats its effective ask as a variable and runs a monotone binary search over an interval
$a_{k}^{\mathrm{eff}}\in[\,\mathrm{ask}_{k},\overline{r}_{k}\,],$ (40)
replaying Phase 2 to check whether $v_{k}$ would still be selected. The search terminates at the largest ask at which $v_{k}$ remains a winner, and the corresponding per-unit reward is set to
$r_{k}^{\mathrm{s}*}=\text{critical ask of }v_{k}.$ (41)
This critical reward is fully determined by others’ reports and system parameters; it does not depend on the actual ask $a_{k}$ as long as $a_{k}\leq r_{k}^{\mathrm{s}*}$. Therefore, increasing $a_{k}$ above $r_{k}^{\mathrm{s}*}$ either has no effect (if $v_{k}$ is still losing) or causes it to lose a profitable contract, while decreasing $a_{k}$ below $r_{k}^{\mathrm{s}*}$ does not increase the reward it receives. As a result, truthful asking maximizes $v_{k}$’s utility, and Off-AIC2 is truthful for AgPs.
Combining the ES-side and AgP-side arguments, we conclude that Off-AIC2 implements a monotone allocation rule with critical-value payments for both buyers and sellers in a single-parameter setting. Hence the mechanism is truthful (dominant-strategy incentive-compatible) for both ESs and AgPs in the offline stage. ∎
Theorem 4.
@@
(i) Phase 1 (pivotal-rank selection). In Phase 1 of Alg. 2, ESs are sorted in nonincreasing order of their average bids $\overline{\mathrm{bid}}_{i}$, and AgPs are sorted in nondecreasing order of their asks $\mathrm{ask}_{k}$. The pivotal ranks $(I^{\ast},K^{\ast})$ are chosen such that
$\overline{\mathrm{bid}}_{I^{\ast}}\;\geq\;\mathrm{ask}_{K^{\ast}},\qquad\sum_{k=1}^{K^{\ast}}C_{k}^{\mathrm{V}}\;\geq\;\sum_{i=1}^{I^{\ast}}N_{i}^{\mathrm{Dem}}.$ (44)
Alg. 2 only considers the top-$I^{\ast}$ ESs and top-$K^{\ast}$ AgPs for matching. Hence, for any winning pair $(i,k)$ with $x^{\mathrm{Off}}_{i,k}=1$, we have
@@
in which both ES $s_{i}$ and AgP $v_{k}$ obtain nonnegative utilities and the auctioneer’s per-unit margin prp-r is nonnegative.
(ii) Phase 3 (critical pricing). Phase 3 of Alg. 2 determines the actual payments within this feasible band using critical-value binary searches:
$\bullet$ Buyer side (ESs): For each winning ES $s_{i}$ with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$, the payment $p_{i}^{\mathrm{b}*}$ is obtained via a monotone binary search on its effective bid parameter. The search interval is chosen inside the buyer side of the surplus-feasible band in (45), and the algorithm stops at the minimum critical bid that keeps $s_{i}$ in the winning set when Phase 2 is re-executed. Thus, for any matched pair $(i,k)$,
@@
$\bullet$ Seller side (AgPs): Similarly, for each winning AgP $v_{k}$ with $\sum_{i}x^{\mathrm{Off}}_{i,k}\geq 1$, the reward $r_{k}^{\mathrm{s}*}$ is obtained via a monotone binary search on its effective ask parameter. The search interval is chosen inside the seller side of the band in (45), and the algorithm stops at the maximum critical ask that keeps $v_{k}$ in the winning set. Hence, for any matched pair $(i,k)$,
@@
while unmatched pairs contribute zero. Summing over all ES–AgP pairs gives
@@
Appendix D Proof Details for the Ordinal Potential and PG-BRD
This appendix collects the main technical ingredients behind the ordinal-potential formulation in Stage II and the design of PG-BRD. We: (i) rewrite the model using a global assignment matrix, (ii) relate pointwise congestion penalties to integrated congestion via a simple sandwich inequality, (iii) characterize single-user deviations and the linearized soft-deadline utility, (iv) prove the ordinal-potential property and finite-improvement path (FIP), and (v) clarify why monetary payments are excluded from the potential.
D-A Matrix Formulation Based on the Global Assignment Matrix
@@
D-B Integral–Pointwise Sandwich Inequality
@@
Lemma 1 (Integral–pointwise sandwich).
@@
D-C Single-User Deviation: Utility and Potential Changes
@@
Payment terms are absent from $\Delta\Phi$ by construction (see Appx. D-I).
D-D Approximate Linearization of Soft-Deadline Utility
We briefly justify the linearized MaU utility in (18) and the definition of the congestion-sensitivity coefficient $\kappa_{\mathbbm{j}}$.
Recall that for MaU $u^{\mathrm{On}}_{\mathbbm{j}}$ offloading to $s^{\mathrm{On}}_{\mathbbm{i}}$,
$G^{\mathrm{A}}_{\mathbbm{j},\mathbbm{i}}=v_{\mathbbm{j}}\,\frac{\big[D_{\mathbbm{j}}-T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}\big]_{+}}{D_{\mathbbm{j}}},\qquad T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}=t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}+t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}.$ (64)
In the practically relevant regime where $T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}\lesssim D_{\mathbbm{j}}$, the operator $\cdot\,$$_{+}$ is active and
$G^{\mathrm{A}}_{\mathbbm{j},\mathbbm{i}}=v_{\mathbbm{j}}-v_{\mathbbm{j}}\frac{t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}}{D_{\mathbbm{j}}}-v_{\mathbbm{j}}\frac{t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}}{D_{\mathbbm{j}}}.$ (65)
@@ D-D Approximate Linearization of Soft-Deadline Utility
$G^{\mathrm{A}}_{\mathbbm{j},\mathbbm{i}}=\underbrace{v_{\mathbbm{j}}-v_{\mathbbm{j}}\frac{t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}}{D_{\mathbbm{j}}}}_{\text{independent of }y_{\mathbbm{i}}}\;-\;\underbrace{v_{\mathbbm{j}}\frac{\theta_{\mathbbm{j},\mathbbm{i}}}{D_{\mathbbm{j}}}\,\tau_{\mathbbm{i}}\big(y_{\mathbbm{i}}\big)}_{\text{congestion-dependent term}}.$ (66)
Adding the energy-saving term $c^{\mathrm{A},\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}=E^{\mathrm{loc}}_{\mathbbm{j}}-E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$ with coefficient $\mathbb{V}^{\mathrm{A}}_{e}$ and subtracting the payment, the full MaU utility can be written as
$U^{\mathrm{A}}_{\mathbbm{j},\mathbbm{i}}=\bar{S}_{\mathbbm{j}}(\mathbbm{i})\;-\;\kappa_{\mathbbm{j}}\,\theta_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}\big(y_{\mathbbm{i}}\big)\;-\;p_{\mathbbm{j},\mathbbm{i}},$ (67)
@@
$\bar{S}_{\mathbbm{j}}(\mathbbm{i})=v_{\mathbbm{j}}-v_{\mathbbm{j}}\frac{t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}}{D_{\mathbbm{j}}}+\mathbb{V}^{\mathrm{A}}_{e}\,c^{\mathrm{A},\mathrm{save}}_{\mathbbm{j},\mathbbm{i}},\qquad\kappa_{\mathbbm{j}}=\frac{v_{\mathbbm{j}}}{D_{\mathbbm{j}}}>0.$ (68)
Thus, MaUs share the same congestion structure as HuUs, with user-specific $\kappa_{\mathbbm{j}}$. This linearization preserves the ordinal ranking of actions induced by the original soft-deadline utility in the operating regime of interest, and hence does not change the set of Nash equilibrium.
D-E Sign Consistency of Congestion Terms
@@
Lemma 2 (Sign consistency).
@@
Apply Lemma 1 to both integrals. Each integral lies between a lower and an upper pointwise congestion term. After scaling by the positive factors $\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}}/\gamma_{\mathbbm{j},\mathbbm{i}}$ and $\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}^{\prime}}/\gamma_{\mathbbm{j},\mathbbm{i}^{\prime}}$, the corresponding bounds match the pointwise expressions in the lemma. Hence the two congestion components must share the same sign. ∎
@@
D-F Monotonicity and Convexity of SP Cost

### v2

@@ B-E Economic Properties of Off-AIC2
Critical-reward-like payment (Step 3). For each winning AP $v_{k}$ (with $\sum_{i}x^{\mathrm{Off}}_{i,k}\geq 1$), Step 3 treats its effective ask as a variable and performs a binary search over
$a_{k}^{\mathrm{eff}}\in[\,\mathrm{ask}_{K^{\ast}+1},\mathrm{ask}_{k}\,],$ (39)
re-running Step 2 to check whether $v_{k}$ would still be selected. The search terminates at the largest ask at which $v_{k}$ remains a winner, and the corresponding reward $r_{k}^{\mathrm{s}*}$ plays the role of a critical ask. When the allocation is exactly monotone, this would give a classical critical payment; with eACO-VRP, $r_{k}^{\mathrm{s}*}$ still behaves as a critical-reward approximation. Empirically (Fig. 2), $v_{k}$’s realized utility as a function of $a_{k}$ exhibits a clear peak near its true ask, and over-reporting beyond that region mostly increases the risk of losing the contract without improving utility. Thus, truthful asking is again a locally best or near-best strategy for APs.
Putting the buyer and seller sides together, Off-AIC2 combines (i) an allocation rule that is monotone up to small perturbations due to the heuristic eACO-VRP subroutine, and (ii) payments and rewards that closely emulate critical values via binary search. According to our definition of approximate truthfulness, this implies that, for both ESs and APs, truthful reporting is a (local) utility-maximizing strategy and any unilateral deviation from the true valuation changes the utility by at most a small margin, as also corroborated by the numerical evidence in Figs. 22. Therefore, Off-AIC2 is near-truthful for both ESs and APs in the offline stage. ∎
Proposition 6.
@@
(i) Phase 1 (pivotal-rank selection). In Step 1 of Alg. 3, ESs are sorted in nonincreasing order of their average bids $\overline{\mathrm{bid}}_{i}$, and APs are sorted in nondecreasing order of their asks $\mathrm{ask}_{k}$. The pivotal ranks $(I^{\ast},K^{\ast})$ are chosen such that
$\overline{\mathrm{bid}}_{I^{\ast}}\;\geq\;\mathrm{ask}_{K^{\ast}},\qquad\sum_{k=1}^{K^{\ast}}C_{k}^{\mathrm{A}}\;\geq\;\sum_{i=1}^{I^{\ast}}N_{i}^{\mathrm{Dem}}.$ (42)
Alg. 3 only considers the top-$I^{\ast}$ ESs and top-$K^{\ast}$ APs for matching. Hence, for any winning pair $(i,k)$ with $x^{\mathrm{Off}}_{i,k}=1$, we have
@@
in which both ES $s_{i}$ and AP $v_{k}$ obtain nonnegative utilities and the auctioneer’s per-unit margin prp-r is nonnegative.
(ii) Phase 3 (critical pricing). Step 3 of Alg. 3 determines the actual payments within this feasible band using critical-value binary searches:
$\bullet$ Buyer side (ESs): For each winning ES $s_{i}$ with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$, the payment $p_{i}^{\mathrm{b}*}$ is obtained via a monotone binary search on its effective bid parameter. The search interval is chosen inside the buyer side of the surplus-feasible band in (43), and the algorithm stops at the minimum critical bid that keeps $s_{i}$ in the winning set when Step 2 is re-executed. Thus, for any matched pair $(i,k)$,
@@
$\bullet$ Seller side (APs): Similarly, for each winning AP $v_{k}$ with $\sum_{i}x^{\mathrm{Off}}_{i,k}\geq 1$, the reward $r_{k}^{\mathrm{s}*}$ is obtained via a monotone binary search on its effective ask parameter. The search interval is chosen inside the seller side of the band in (43), and the algorithm stops at the maximum critical ask that keeps $v_{k}$ in the winning set. Hence, for any matched pair $(i,k)$,
@@
while unmatched pairs contribute zero. Summing over all ES–AP pairs gives
@@
Appendix C Properties of the Ordinal Potential
To turn the above game formulation into a practical distributed scheduling mechanism, we need a global scalar function whose improvement is aligned with individual SD incentives. In particular, if $\Phi(\bm{\pi})$ acts as an ordinal potential, then any unilateral utility-improving move by an SD strictly increases $\Phi(\bm{\pi})$. This property is exactly what allows simple asynchronous better-/best-response dynamics to converge to a pure-strategy NE with provable guarantees and low online complexity. We therefore establish that the function $\Phi(\bm{\pi})$ in (21) is an ordinal potential for the online game under soft-deadline valuations.
The proof follows a standard single-user deviation argument and relies on three ingredients: (i) separability of the non-congestion terms $S_{\mathbbm{j}}(\cdot)$, (ii) an integral–pointwise comparison that links the pointwise congestion penalty in user utilities with the integrated congestion cost in $\Phi(\bm{\pi})$, and (iii) monotonicity and convexity of the SP-side cost $K_{\mathbbm{i}}(\cdot)$. Full details are deferred to Appendices E-CE-G.
Proposition 7 (Ordinal potential).
Consider the unified utilities in (19) with soft-deadline MUs and the potential $\Phi(\bm{\pi})$ in (21). For any SD $u^{\mathrm{On}}_{\mathbbm{j}}$ and any unilateral deviation $\pi_{\mathbbm{j}}=\mathbbm{i}\to\mathbbm{i}^{\prime}$, if $U_{\mathbbm{j},\mathbbm{i}^{\prime}}-U_{\mathbbm{j},\mathbbm{i}}>0$, then $\Phi\big(\bm{\pi}_{-\mathbbm{j}},\mathbbm{i}^{\prime}\big)-\Phi\big(\bm{\pi}_{-\mathbbm{j}},\mathbbm{i}\big)>0$. Hence, $\Phi(\bm{\pi})$ is an ordinal potential for the considered game[13].
Since each SD possesses a finite action set, any sequence of unilateral best-response (or better-response) updates generates a strictly increasing trajectory of the potential function $\Phi(\bm{\pi})$ and therefore converges in finitely many iterations to a pure-strategy NE. Equivalently, every local maximizer of $\Phi(\bm{\pi})$ corresponds to a pure-strategy NE of the online scheduling game. A formal proof of this finite-improvement property is provided in Appx. E-H.
Appendix D Detailed pseudo-code of PG-BRD
In this section, we provide the detailed pseudo-code of PG-BRD in Alg. 4.
1Input: SD set $\mathcal{U}^{\mathrm{On}}$, SP set $\mathcal{S}^{\mathrm{On}}$, coverage sets $\{\bm{C}_{\mathbbm{j}}\}$, subcarrier-capacity $\{G_{\mathbbm{i}}\}$, max iterations $T_{\max}$, threshold $\varepsilon$.
2[2pt]
3Initialization: Construct a random feasible profile $\bm{\pi}=(\pi_{\mathbbm{j}})_{\mathbbm{j}\in\mathcal{U}^{\mathrm{On}}}$ that satisfies constraints (22a)–(22c); set $\mathbbm{t}\leftarrow 0$ and $\mathrm{converged}\leftarrow\textbf{false}$.
4[2pt]
5while not $\mathrm{converged}$ and $\mathbbm{t}<T_{\max}$ do
6
7 $\mathrm{converged}\leftarrow\textbf{true}$;
8  Generate a random permutation $\mathcal{P}$ of indices in $\mathcal{U}^{\mathrm{On}}$;
9
10 for each SD $u^{\mathrm{On}}_{\mathbbm{j}}$ in $\mathcal{P}$ do
11     let $a^{\mathrm{cur}}_{\mathbbm{j}}\leftarrow\pi_{\mathbbm{j}}$;
12     for all $\mathbbm{i}$, compute the loads excluding $u^{\mathrm{On}}_{\mathbbm{j}}$: $y^{-{\mathbbm{j}}}_{\mathbbm{i}}$, $\mathbbm{x}^{-{\mathbbm{j}}}_{\mathbbm{i}}$, and $n^{-{\mathbbm{j}}}_{\mathbbm{i}}$;
13
14    build the feasible action set $\mathcal{A}^{\mathrm{feas}}_{\mathbbm{j}}$ according to (23);
15
16    compute current utility $U_{\mathbbm{j}}(a^{\mathrm{cur}}_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})$ via (19);
17
18    for each $a\in\mathcal{A}^{\mathrm{feas}}_{\mathbbm{j}}$ do
19        temporarily set $\pi_{\mathbbm{j}}\leftarrow a$ and evaluate $U_{\mathbbm{j}}(a\mid\bm{\pi}_{-\mathbbm{j}})$;
20
21
22    let $a^{\mathrm{best}}_{\mathbbm{j}}\in\arg\max_{a\in\mathcal{A}^{\mathrm{feas}}_{\mathbbm{j}}}U_{\mathbbm{j}}(a\mid\bm{\pi}_{-\mathbbm{j}})$;
23
24    if $U_{\mathbbm{j}}(a^{\mathrm{best}}_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})-U_{\mathbbm{j}}(a^{\mathrm{cur}}_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})>\varepsilon$ then
25        $\pi_{\mathbbm{j}}\leftarrow a^{\mathrm{best}}_{\mathbbm{j}}$;
26        $\mathrm{converged}\leftarrow\textbf{false}$;
27    else
28        $\pi_{\mathbbm{j}}\leftarrow a^{\mathrm{cur}}_{\mathbbm{j}}$;
29
30
31 $\mathbbm{t}\leftarrow\mathbbm{t}+1$;
32
Output: NE profile $\bm{\pi}^{\star}=\bm{\pi}$ and the associated allocation.
Algorithm 4 Proposed PG-BRD
Appendix E Proof Details for the Ordinal Potential and PG-BRD
This appendix collects the main technical ingredients behind the ordinal-potential formulation in the online stage and the design of PG-BRD. We have: (i) rewrite the model using a global assignment matrix, (ii) relate pointwise congestion penalties to integrated congestion via a simple sandwich inequality, (iii) characterize single-user deviations and the linearized soft-deadline utility, (iv) prove the ordinal-potential property and finite-improvement path (FIP), and (v) clarify why monetary payments are excluded from the potential.
E-A Matrix Formulation Based on the Global Assignment Matrix
@@
E-B Integral-Pointwise Sandwich Inequality
@@
Proposition 8 (Integral-pointwise sandwich).
@@
E-C Single-User Deviation: Utility and Potential Changes
@@
Payment terms are absent from $\Delta\Phi$ by construction (see Appx. E-I).
E-D Approximate Linearization of Soft-Deadline Utility
We briefly justify the linearized MU utility in (18) and the definition of the congestion-sensitivity coefficient $\kappa_{\mathbbm{j}}$.
Recall that for MU $u^{\mathrm{On}}_{\mathbbm{j}}$ offloading to $s^{\mathrm{On}}_{\mathbbm{i}}$,
$G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=v_{\mathbbm{j}}\,\frac{\big[\tau^{\mathrm{ddl}}_{\mathbbm{j}}-T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}\big]_{+}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}},\qquad T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}=t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}+t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}.$ (62)
In the practically relevant regime where $T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}\lesssim\tau^{\mathrm{ddl}}_{\mathbbm{j}}$, the operator $\cdot\,$$_{+}$ is active and
$G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=v_{\mathbbm{j}}-v_{\mathbbm{j}}\frac{t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}}-v_{\mathbbm{j}}\frac{t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}}.$ (63)
@@ E-D Approximate Linearization of Soft-Deadline Utility
$G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=\underbrace{v_{\mathbbm{j}}-v_{\mathbbm{j}}\frac{t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}}}_{\text{independent of }y_{\mathbbm{i}}}\;-\;\underbrace{v_{\mathbbm{j}}\frac{\theta_{\mathbbm{j},\mathbbm{i}}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}}\,\tau_{\mathbbm{i}}\big(y_{\mathbbm{i}}\big)}_{\text{congestion-dependent term}}.$ (64)
Adding the energy-saving term $c^{\mathrm{M},\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}=E^{\mathrm{loc}}_{\mathbbm{j}}-E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$ with coefficient $\mathbb{V}^{\mathrm{M}}_{e}$ and subtracting the payment, the full MU utility can be written as
$U^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=\bar{S}_{\mathbbm{j}}(\mathbbm{i})\;-\;\kappa_{\mathbbm{j}}\,\theta_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}\big(y_{\mathbbm{i}}\big)\;-\;p_{\mathbbm{j},\mathbbm{i}},$ (65)
@@
$\bar{S}_{\mathbbm{j}}(\mathbbm{i})=v_{\mathbbm{j}}-v_{\mathbbm{j}}\frac{t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}}+\mathbb{V}^{\mathrm{M}}_{e}\,c^{\mathrm{M},\mathrm{save}}_{\mathbbm{j},\mathbbm{i}},\qquad\kappa_{\mathbbm{j}}=\frac{v_{\mathbbm{j}}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}}>0.$ (66)
Thus, MUs share the same congestion structure as HUs, with user-specific $\kappa_{\mathbbm{j}}$. This linearization preserves the ordinal ranking of actions induced by the original soft-deadline utility in the operating regime of interest, and hence does not change the set of NE.
E-E Sign Consistency of Congestion Terms
@@
Proposition 9 (Sign consistency).
@@
Apply Proposition 8 to both integrals. Each integral lies between a lower and an upper pointwise congestion term. After scaling by the positive factors $\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}}/\gamma_{\mathbbm{j},\mathbbm{i}}$ and $\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}^{\prime}}/\gamma_{\mathbbm{j},\mathbbm{i}^{\prime}}$, the corresponding bounds match the pointwise expressions in the lemma. Hence the two congestion components must share the same sign. ∎
@@
E-F Monotonicity and Convexity of SP Cost